The sum of a number and -16 is 3. Find the number.
Found 2 solutions by Ayesha, jim_thompson5910:
Answer by Ayesha(2)

Let the number be x. Thus x-16=3 this implies x=16+3 Therefore x=18

Answer by jim_thompson5910(35256)

Correction:
x=16+3
x=19
Just fixing a minor typo